OpenStudy (riot2015):

I NEED HELP!I am a 15 year old girl with a 76 beats per minute and im trying to find out % target heart rate i should have.

OpenStudy (jebonna):

Work out your maximum heart rate: 220 - 15 = 205bpm This is usually used for exercising, so it will give you the range of heart rate in bpm of what you should have or aim for whilst exercising: 205 / 100 x 50 = 103bpm 205 / 100 x 70 = 144bpm So when exercising you should aim for a range of bpm of 103bpm-144bpm The average heart rate whilst resting is 70bpm-100bpm

OpenStudy (riot2015):

so im supposed to answer these: Maximum Heart Rate: (220-age): Target Heart Rate: (220 – age x .65 to .85) : ....so i should put 205 for the first one and what for the second one?

OpenStudy (riot2015):

@jebonna

OpenStudy (riot2015):

I dont get how i am supposed to pic a number between .65 and .85